OpenStudy (anonymous):

find the inverse function of 4,7 6,2 9,-1 11,3

OpenStudy (anonymous):

so 4 goes to 7 ( add 3) --> f(x) = x+3 the inverse function is f(x) = 1/x+3 so 6 goes to 2 (subtract 4) ---> f(x) = x - 4 so inverse function f(x) = 1/x-4 You get the idea!